OpenStudy (anonymous):

rewrite logarithm using common logs. then evaluate using a calculator... i missed class this day so i dont know how to do this..

OpenStudy (anonymous):

\[\log _{3}0.25\]

OpenStudy (anonymous):

how exactly do you want me to rewrite the logarithm??

OpenStudy (anonymous):

it just says using common logs

OpenStudy (anonymous):

Well, I am not sure what you exactly want. If you give me any example or the answer you are looking for, then maybe I can help you

OpenStudy (anonymous):

well can u try this one... write the logarithmic expression in expanded form \[\log _{6}x ^{2}/y ^{3}\]

OpenStudy (anonymous):

Sure Use these properties and get your answer \[\log_{a}b ^{c} = c*\log_{a}b \] and this \[\log_{a}(b/c) = \log_{a}(b)- \log_{a}(c)\] Tip: apply 2nd property first and then the first one

OpenStudy (precal):